### v1

# Enhanced Quantum Energy Teleportation using a 3-Qubit System

Abstract
Quantum Energy Teleportation (QET) is a novel method that leverages quantum entanglement to transfer energy between two distant locations without any physical movement of the energy. The first realization of QET on superconducting hardware, utilizing a 2-qubit system, demonstrated an average energy retrieval efficiency of 11.4% by the receiver, Bob. In this paper, we present a new approach using a 3-qubit system to enhance the energy efficiency of QET. We have incorporated a novel 3-qubit ground state hamiltonian $H$ to achieve this, that conforms the constraints of Zero mean energy. Our experimental results show a significant improvement in energy retrieval, achieving an average efficiency of 46.4%, which is significantly higher than that of the 2-qubit system. This advancement not only marks a step forward in practical quantum energy applications but also provides a new framework for future research in quantum energy teleportation and related quantum technologies.
Index Terms:
Quantum Energy Teleportation, QET, 3-Qubit System, Quantum Entanglement, Energy Efficiency.
I Quantum Energy Teleportation (QET)
Alongside the fact that information about quantum state teleportation to faraway locations is widely recognised [1, 2, 3, 4], it is also as widely understood that quantum state energy can also be transmitted in a similar manner, paves the potential for future uses. Quantum information transmitted through quantum teleportation is not a tangible measure, whereas energy is a clearly defined physical measure. Transmitting physical quantities to distant locations is a somewhat uncharted domain of technology. Quantum Energy Teleportation (QET) was first suggested by Hotta approximately 17 years ago. Since then, it was the subject of theoretical investigation in spin chains [5, 6, 7], a quantum Hall system [9], an ion trap system[8] and other diverse systems [10, 11] taht are still at a theoretical level. It is remarkable that the experimental validation of QET has been infrequent prior to [33], despite its achievability using a straightforward quantum system. The initial empirical validation of Quantum Energy Teleportation (QET) using real cloud-based quantum computers has been conducted in [34] in a very prominent manner, with necessary quantum circuits to do this. They successfully implemented Quantum Error Correction (QET) on IBM quantum environment that leverages superconducting quantum computers by employing quantum error mitigation techniques [13, 14, 15] as well.
The purpose of this paper is to ameliorate the efficiency of the energy retrieved from this protocol by extending the number of qubits used.The quantum hardware utilised in our study is IBM’s quantum computer, ibm_fez, which is accessible to everyone worldwide at no cost. By utilising the quantum circuits presented in this research, individuals will have the capability to replicate the outcomes and quantum energy teleportation (QETs) in an efficient manner. Given that the features of quantum computers are openly accessible nowadays, it will be feasible for everyone to use the extended QET protocol. The techniques we have introduced on top of minimal qet model [34] can be utilised on any system that has the capability of Quantum Energy Teleportation (QET).
The following explanation clarifies why QET serves as a universal method for quantum energy transfer, similar to how quantum teleportation serves the role of universal method for energy and quantum information transfer. Excited states are indigenous to the observations of the ground state of a quantum many-body system, which subsequently raise the expected energy level. It should be noted that the experimental devices provide the additional energy. The ground state of a quantum many-body system possesses the significant characteristic of entanglement, which results in quantum fluctuations in the overall energy of the ground state. To clarify, the energy fluctuations of the local systems are entangled due to quantum effects. Performing a measurement on subsystem A at the local level results in the destruction of the entanglement of the ground state. Simultaneously, the measurement instrument transfers energy $E_{A}$ into the entire system. The injected energy $E_{A}$ remains localised within subsystem A throughout the initial phases of time evolution. However, activities solely focused on subsystem A are unable to withdraw $E_{A}$ from the system. This is because information regarding $E_{A}$ is distributed over remote places in addition to A, as a result of the pre-existing entanglement. Put simply, the energy injected locally, denoted as $E_{A}$, can be partially retrieved from any point except A [16]. The QET protocol enables this capability. Thus yet, no specific assumptions have been made regarding the system. The key characteristic of QET is its complete realisation through the inherent properties of the quantum many-body system ground state and the universally observed phenomenon called measurement.
The QET model described in [17] is a simple implementation that employs real quantum networks and quantum computers in a quantum circuit. However, the limitation resides in the efficiency of the teleportation process, as BoB is only able to recover approximately 11% of the energy that Alice initially produced. This research utilised an expanded iteration of the aforementioned method, employing quantum circuits consisting of three qubits for Quantum Error Correction (QET), as illustrated in Figure 1. Quantum computers already possess sufficient capability to execute a circuit depth of up to 9.
(a) Quantum Energy Teleportation (Minimal)
(b) Quantum Energy Teleportation (Extended)
Figure 1: Comparison of Quantum Energy Teleportation
I-A Defining The Essentials Of QET
To begin, we will provide a comprehensive overview of the QET protocol[37]. To find quantum circuit implementations for specific situations, refer to [33, 34, 35, 36] and Fig. 2. where local Hamiltonian as $H=\sum_{n=0}^{N}H_{n}$, is defined and here $H_{n}$ represents the local Hamiltonian that interacts with surrounding qubits. It must satisfy the following constraints
$\langle g|H|g\rangle=\langle g|H_{n}|g\rangle=0,\quad\forall n\in\{1,\cdots,N\},$ (1)
Here $|g\rangle$ is said to be the ground state of the total Hamiltonian $H$. But in case of local $H_{n}$ it might always not be the case. It is crucial to acknowledge that $|g\rangle$ is a state of entanglement in a generic context. To uphold the requirement (1), it is possible to consistently sum or deduct constant values, As the ground state, it is evident that any non-trivial (local) operations to $|g\rangle$ such as measurement, results in increase of the energy expectation value.
Below, we provide a description of the QET protocol. Alice plays the role of energy supplier while Bob remains as a receiver. Alice does a projective measurement on her Pauli operator $\sigma_{A}$, using operator $P_{A}(\mu)=\frac{1}{2}(1+\mu\sigma_{A})$. The results she obtains is either $\mu=-1$ or $\mu=+1$. $E_{A}$ that is the injected energy is localized around subsystem $A$, but Alice is unable to withdraw it from the system merely through her operations at $A$. Nevertheless, by employing LOCC, Bob has the ability to extract a certain amount of energy from his local system.
Alice transmits her measurement result $\mu$ to Bob by classical communication. Upon receiving the result, Bob applies conditional operation $U_{B}(\mu)$ to his state and perfroms a measurement on his local Hamiltonian $H_{B}$. His operation can be defined as
$U_{B}(\mu)=\cos\theta I-i\mu\sin\theta\sigma_{B},$ (2)
where $\theta$ obeys
$\cos(2\theta)=\frac{\xi}{\sqrt{\xi^{2}+\eth^{2}}},\sin(2\theta)=-\frac{\eta}{\sqrt{\xi^{2}+\eth^{2}}}$ (3)
where
$\xi=\langle g|\sigma_{B}H\sigma_{B}|g\rangle,\quad\eta=\langle g|\sigma_{A}\sigma_{B}|g\rangle,$ (4)
Here $\dot{\sigma}_{B}=i[H,\sigma_{B}]$. $[H,\sigma_{B}]=[H_{B},\sigma_{B}]$ must be maintained by the local hamiltonian. The average quantum state $\rho_{\text{QET}}$, additionally a mixed state, can be gained after applying Bob’s operator $U_{B}(\mu)$ to $\frac{1}{\sqrt{p(\mu)}}P_{A}(\mu)|g\rangle$, where $p(\mu)$ is depicted as a normalization factor.
We get the density matrix $\rho_{\text{QET}}$ after Bob applies the operator $U_{B}(\mu)$ to $P_{A}(\mu)|g\rangle$ is
$\rho_{\text{QET}}=\sum_{\mu\in\{\pm 1\}}U_{B}(\mu)P_{A}(\mu)|g\rangle\langle g|P_{A}(\mu)U_{B}^{\dagger}(\mu).$ (5)
Bob’s expected energy at his local system can be measured as
$\langle E_{B}\rangle=\text{Tr}[\rho_{\text{QET}}H_{B}]=\frac{1}{2}\left[\xi-\sqrt{\xi^{2}+\eth^{2}}\right],$ (6)
which is evaluated negative if $\eta\neq 0$ and no energy dissipation, the positive energy of $-\langle E_{B}\rangle$ is teleported to Bobs device by the law of energy conservation.
I-B Minimal QET Model
For the full description of minimal model refer to [34] Let’s assume $k,h$ be positive numbers. The minimal model is defined as
$H_{\text{tot}}$ $=H_{0}+H_{1}+V,$ (7)
$H_{n}$ $=hZ_{n}+\frac{h^{2}}{\sqrt{h^{2}+k^{2}}},\quad(n=0,1)$ (8)
$V$ $=2kX_{0}X_{1}+\frac{2k^{2}}{\sqrt{h^{2}+k^{2}}}.$ (9)
The ground state of $H_{\text{tot}}$ can be defined as
$|g\rangle=\frac{1}{\sqrt{2}}\sqrt{1-\frac{h}{\sqrt{h^{2}+k^{2}}}}|00\rangle-\frac{1}{\sqrt{2}}\sqrt{1+\frac{h}{\sqrt{h^{2}+k^{2}}}}|11\rangle,$ (10)
One can add the constant terms to the Hamiltonians so that the ground state $|g\rangle$ of $H_{\text{tot}}$ evaluates as zero mean energy for all local and global Hamiltonians:
$\langle g|H_{\text{tot}}|g\rangle=\langle g|H_{0}|g\rangle=\langle g|H_{1}|g\rangle=\langle g|V|g\rangle=0.$ (11)
As we discussed earlier, $|g\rangle$ is cannot be a ground state nor an eigenstate of $H_{n},V,H_{n}+V\,(n=0,1)$. The primary objective seems be obtaining negative ground state energy of local and semi-local Hamiltonians from QET protocol.
Figure 2: Implementation of Minimal QET Model
The QET protocol is outlined below. Alice initially performs a measurement on her Pauli operator $X_{0}$ by $P_{0}(\mu)=\frac{1}{2}(1+\mu X_{0})$ resulting in $\mu=-1$ or $+1$. t Alice’s expectation energy is denoted as,
$E_{0}=-\frac{h^{2}}{\sqrt{h^{2}+k^{2}}}.$ (12)
She communicates her measurement result $\mu$ to Bob using a classical channel, who conducts an operation $U_{1}(\mu)$ to his qubit and measures $H_{1}$ and $V$. The density matrix $\rho_{\text{QET}}$ is evaluated after Bob operates $U_{1}(\mu)$ to $P_{0}(\mu)|g\rangle$ as
$\rho_{\text{QET}}=\sum_{\mu\epsilon\{-1,1\}}U_{1}(\mu)P_{0}(\mu)|g\rangle\langle g|P_{0}(\mu)U_{1}^{\dagger}(\mu).$ (13)
By using $\rho_{\text{QET}}$, the expected local energy at Bob’s subsystem is calculated as $\langle E_{1}\rangle=\text{Tr}[\rho_{\text{QET}}(H_{1}+V)]$, which comes out negative in general. By the law of energy conservation, $E_{B}=-\langle E_{1}\rangle(>0)$ is extracted from the system by the device that operates $U_{1}(\mu)$ [18].
II Quantum Circuit Implementation of Extended QET Model
We define a novel 3-qubit hamiltonian $H_{\text{tot}}$ , where $h\&k$ are positive integers as like minimal qubit model.
{strip}
II-A Defining 3 qubit hamiltonian
$H_{\text{tot}}$ $=H_{0}+H_{1}+H_{2}+V_{0,1}+V_{1,2}+V_{0,2},$ (14)
$H_{n}$ $=hZ_{n}+\frac{x}{\sqrt{h^{2}+hk+k^{2}}+h-k},\quad(n=0,1,2)$ (15)
$V_{i,j}$ $=2kX_{i}X_{j}+\frac{5h^{2}+2hk+5k^{2}+4(h-k)\sqrt{h^{2}+hk+k^{2}}+2k^{2}-x}{\sqrt{h^{2}+hk+k^{2}}+h-k}.$ (16)
where $x$ is defined by
$-a+\left(-a+L\right)\left(M_{1}\right)^{2}\quad+\left(a+L\right)\left(M_{2}\right)^{2}\quad+\left(a+L\right)\left(M_{3}\right)^{2}\quad+L=0$ (18)
Here (19)
$K=\sqrt{h^{2}+hk+k^{2}},$ (20)
$L=\frac{5h^{2}+2hk+4hK+5k^{2}-4cK-x}{3(h-k)+6K},$ (21)
$M_{1}=\frac{8h^{3}k-4h^{2}k^{2}+8h^{2}kK+5hk^{3}-8ak^{2}K-6k^{4}+6k^{3}K}{32h^{4}+32h^{3}K+18h^{2}k^{2}-16h^{2}kK-11hk^{3}+14ak^{2}K+6k^{4}-6k^{3}K},$ (22)
$M_{2}=\frac{32h^{5}k-16h^{4}k^{2}+32h^{4}kK+30h^{3}k^{3}-32h^{3}k^{2}K-26h^{2}k^{4}+34h^{2}k^{3}K+19hk^{5}-25hk^{4}K-12k^{6}+12k^{5}K}{128h^{6}+128h^{5}K+112h^{4}k^{2}-64h^{4}kK-68h^{3}k^{3}+96h^{3}k^{2}K+54h^{2}k^{4}-68h^{2}k^{3}K-31hk^{5}+37hk^{4}K+12k^{6}-12k^{5}K},$ (23)
$M_{3}=\frac{128h^{6}k-64h^{5}k^{2}+128h^{5}kK+144h^{4}k^{3}-128h^{4}k^{2}K-128h^{3}k^{4}+160h^{3}k^{3}K+106h^{2}k^{5}-136h^{2}k^{4}K-69hk^{6}+87hk^{5}K+36k^{7}-36k^{6}K}{512h^{7}+512h^{6}K+544h^{5}k^{2}-256h^{5}kK-320h^{4}k^{3}+480h^{4}k^{2}K+306h^{3}k^{4}-368h^{3}k^{3}K-202h^{2}k^{5}+250h^{2}k^{4}K+105hk^{6}-123hk^{5}K-36k^{7}+36k^{6}K}$
The ground state of $H_{\text{tot}}$ is
$|g\rangle=\mathbf{C}|\psi\rangle^{\dagger},$ (24)
where $C$ is defined as -
$\begin{pmatrix}0&0&0&0&0&0&0&0\\[10.0pt] 0&-M_{3}&0&0&0&0&0&0\\[10.0pt] 0&0&-M_{2}&0&0&0&0&0\\[10.0pt] 0&0&0&0&0&0&0&0\\[10.0pt] 0&0&0&0&-M_{1}&0&0&0\\[10.0pt] 0&0&0&0&0&0&0&0\\[10.0pt] 0&0&0&0&0&0&0&0\\[10.0pt] 0&0&0&0&0&0&0&1\end{pmatrix}$
Here
$K=\sqrt{h^{2}+hk+k^{2}},$
[…]
$M_{2}=\frac{32h^{5}k-16h^{4}k^{2}+32h^{4}kK+30h^{3}k^{3}-32h^{3}k^{2}K-26h^{2}k^{4}+34h^{2}k^{3}K+19hk^{5}-25hk^{4}K-12k^{6}+12k^{5}K}{128h^{6}+128h^{5}K+112h^{4}k^{2}-64h^{4}kK-68h^{3}k^{3}+96h^{3}k^{2}K+54h^{2}k^{4}-68h^{2}k^{3}K-31hk^{5}+37hk^{4}K+12k^{6}-12k^{5}K},$
$M_{1}=\frac{8h^{3}k-4h^{2}k^{2}+8h^{2}kK+5hk^{3}-8hk^{2}K-6k^{4}+6k^{3}K}{32h^{4}+32h^{3}K+18h^{2}k^{2}-16h^{2}kK-11hk^{3}+14hk^{2}K+6k^{4}-6k^{3}K}$
and $|\psi\rangle$ is defined as
$|\psi\rangle=|000\rangle+|001\rangle+|010\rangle+|011\rangle+|100\rangle+|101\rangle+|110\rangle+|111\rangle$ (25)
Figure 3: Implementation of Extended QET Model
{strip}
II-B Deposit Energy (Alice)
Alice uses the below projective measurement operator
$P_{0}(\mu_{0})=\frac{1}{2}(1+\mu_{0}X_{0}).$ (26)
Measurement of Alice’s $X$ operator is conducted, by which we can witness a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. Bob’s energy is certainly not affected by this operation since $[X_{0},V_{0,1}]=[X_{0},H_{1}]=0$. Using $[P_{0}(\mu_{0}),V_{0,1}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, Alice’s mean energy to deposit can be calculated as,
$\langle E_{0}\rangle=\sum_{\mu\in\{-1,1\}}\langle g|P_{0}(\mu_{0})H_{\text{tot}}P_{0}(\mu_{0})|g\rangle$ (27)
II-C Deposit Energy (Charlie)
Now, Charlie uses the following projective measurement operator
$P_{1}(\mu_{1})=\frac{1}{2}(1+\mu_{1}X_{1}).$ (28)
Charlie’s $X$ operator is measured, by which he obtains a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. His operation also does not affect Bob’s energy since $[X_{1},V_{0,2}]=[X_{0},H_{2}]=0$. Using $[P_{0}(\mu),V_{0,2}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, we find that Charlie’s mean energy to deposit is
$\langle E_{1}\rangle=\sum_{\mu\in\{-1,1\}}\langle g|P_{0}(\mu_{0})|P_{1}(\mu_{1})H_{\text{tot}}|P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle$ (29)
Step 2: Receive Energy
As soon as Alice observes $\mu_{0}\in\{-1,1\}$ and Chalie observes $\mu_{1}\in\{-1,1\}$ , They communicates their result to Bob who operates $U_{B}(\mu_{0,1})$ to his qubit and measures his energy. Here $U_{B}(\mu_{0,1})$ is obtained as
$U_{1}(\mu_{0,1})=\cos\phi I-i\mu\sin\phi Y_{1}=R_{Y}(2\phi),$ (30)
The average quantum state obtained after Bob operates $U_{B}(\mu)$ to $P_{0}(\mu)|g\rangle$ is
$\rho_{\text{QET}}=\sum_{\mu\in\{-1,1\}}U_{1}(\mu_{0,1})P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu_{0,1})$ (31)
Then the average energy measured by Bob is
$\langle E_{2}\rangle=\text{Tr}[\rho_{\text{QET}}(H_{2}+V_{0,1}+V_{0,2})]=\text{Tr}[\rho_{\text{QET}}H_{\text{tot}}]-(\langle E_{0}\rangle+\langle E_{1}\rangle),$ (32)
$[U_{1}(\mu),H_{1}]=0$ is beign employed here. It is evident that the mapping of $\sum_{\mu\in\{-1,1\}}P_{0}(\mu)[\langle g|P_{0}(\mu)\rangle]\rightarrow\rho_{\text{QET}}$ is not a unitary transformation. Thus, in contrast to eq. (27) and (29) , eq. (32) can be negative.
Here is a little bit explanation of quantum circuits for the extended QET protocol. Similar to minimal QET model $V$ and $H_{2}$ do not commute, measurement of those has to be done separately. That means, Bob measures $V_{0,2}$,$V_{0,1}$ and $H_{2}$ independently and obtains evaluated $V_{0,2}$,$V_{0,1}$ and $\langle H_{2}\rangle$ statistically. We witness, $V_{0,2}$,$V_{1,2}$ to be always negative and $\langle H_{2}\rangle$ is always positive. Therefore it is sufficient for Bob to measure only $V_{0,2}$,$V_{1,2}$ to receive energy with our extended QET protocol.
We can consider $V(\mu)=\langle g|P_{0}(\mu)[U_{1}(\mu)VU_{1}^{\dagger}(\mu)]P_{0}(\mu)|g\rangle$. Fig. 1 shows the quantum circuit to measure $V(\mu)$ which on the right panel of Fig. 1 (B). It is important to note that, although $V$ is not a local operator, since Bob’s measurement is dependent on Alice’s data as well as Chalie’s, we obtain $\rho_{\text{QET}}(\mu)$ by Bobs’s local measurement only. Similarly $H_{2}$ can be measured as well as shown in the left panel of Fig. 3 (A) with a Z-basis gate. The average the energy expectation value generated by our curcuit is,
$\langle E_{2}\rangle=\sum_{\mu\in\{-1,1\}}\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu)(H_{2}+V_{0,1}+V_{0,2})U_{1}(\mu)P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle.$ (33)
$\phi$ is tested with different real values and we have noticed negative $\langle E_{2}\rangle$ value as expected. Bob receives energy $\langle E_{B}\rangle=-\langle E_{2}\rangle$ on average.
III Extended QET on Real Quantum Computer
we explain the process of executing conditional operations on real quantum hardware, specifically focusing on operations that are not inherently supported by most quantum computers and devices. In the QET protocol, Bob’s operation must be chosen based on the outcomes of Alice’s measurements, as seen in Figure. 3 (B) is a notation used to indicate the second item or option in a list or sequence.
QET can be applied seamlessly in settings that do not support conditional statements by utilizing the deferred measurement technique. By deferring Alice’s measurement to the conclusion of the circuit, we can achieve identical outcomes. The conditional operations can be constructed using a controlled $U$ gate $\Lambda(U)=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes U$. One would find the equivalence between the following two circuits enclosed by the orange dashed frame in Fig. 2.
Utilization of the circuit enclosed by the orange dashed frame are in Figure.3.In case of extended model we have conducted quantum computation utilizing 3 distinct IBM quantum hardware devices: ibm_fez, ibm_torino, ibmq_hanoi. The characteristics of each quantum computer can be observed in later Figures. A CNOT gate can be directly applied to two qubits that are coupled at the edge. To do a quantum computation, we can select two qubits that are positioned on the connected edges. The experiment was conducted by measuring the predicted values, which exhibited a high degree of similarity across different devices.
Ultimately, we conducted a simulation utilizing a qasm simulator, which has the capability to accurately replicate operations done on the identical quantum circuits employed for quantum computation. In the following part, we will provide a concise overview of the outcomes achieved using quantum computation. The findings obtained from the simulator were in agreement with the analytical solution, demonstrating a high level of precision. This confirms that the quantum circuit was built appropriately. Additional experimental findings are consolidated in Table IV located in the Appendix. The specific characteristics of the machine and the experimental parameters are outlined later.
(a) Bob Measures $V_{0,1}$ $V_{0,2}$
(b) Bob Measures $H_{2}$
In this study we discovered negative energy, denoted $\langle E_{1}\rangle<0$ and the value is smaller that what we can obtain in case of 2-qubit QET , which can be considered a new . The value of $\langle V\rangle$ that was closest to the exact analysis value was -4.4507 ($h=18.24,k=7.15$ with ibm_torino), which is about 86% accurate. As emphasized in Hotta’s original works [5, 6, 7, 8, 9, 10, 11, 17],it is impossible for any unitary operation to cause $\langle E_{1}\rangle$ to become negative after Alice observes $X_{0}$(eq. 21). In order for Bob to get the accurate $\langle E_{1}\rangle$, Alice, Chalie and Bob must conduct the experiment a vast number of times, and the corresponding value of $\langle V\rangle$ and $\langle H_{2}\rangle$ can be obtained only when Alice, Chalie and Bob communicate appropriately in the quantum circuit in Fig. 1 (B). Distributions of states obtained by a quantum computer ibm_kyiv are shown in Fig. 4 (B), where distributions of raw results are shown with a simulator qasm_simulator. In this study, we compare the distributions of raw findings and error-mitigated outcomes from 2-qubit system vs 3-qubit system using a simulator called qasmsimulator.
We employed a straightforward measurement error reduction technique to assess the impact of measurement mistakes. We compiled a roster of four measurement calibration circuits designed for the complete Hilbert space. Subsequently, we promptly conducted measurements in order to acquire the probability distributions. Subsequently, we utilized the calibration matrix to rectify the obtained measurements. The mean measurement fidelity for each quantum computer is presented later. The histograms of the observed states exhibited consistent patterns across all other quantum computers we utilized. The histograms generated from the measurement of $H_{2}$ exhibit a high level of agreement with the simulator results, demonstrating good accuracy. The enhancement of the values resulting from the avoidance of measurement errors is also verified by the outcomes presented in Table I. The observation of visibility is of paramount significance in this investigation. While the raw data obtained from quantum computers showed discrepancies compared to the results from the simulator, error mitigation techniques were able to enhance the data to the extent that negative energy expectation values
Backend 2 Qubit QET 3 Qubit QET
[…]
$(E_{0})$
Analytical value .7071 1.2481 5.23 4.67
qasm_simulator 0.7088 $\pm$ 0.0001 1.2437 $\pm$ 0.0047 5.23 $\pm$ 0.0047 4.67 $\pm$ 0.0047
ibm_fez error mitigated 0.6560 $\pm$ 0.0031 1.2480 $\pm$ 0.0047 5.23 $\pm$ 0.0056 4.65 $\pm$ 0.0027
unmitigated 0.6874 $\pm$ 0.0031 1.4066 $\pm$ 0.0047 5.24 $\pm$ 0.0017 4.67 $\pm$ 0.0024
ibm_torino error mitigated 0.7039 $\pm$ 0.0056 1.2318 $\pm$ 0.0084 5.23 $\pm$ 0.0054 4.67 $\pm$ 0.0078
unmitigated 0.7232 $\pm$ 0.0056 1.2624 $\pm$ 0.0083 5.24 $\pm$ 0.0053 4.66 $\pm$ 0.0033
ibm_kyiv error mitigated 0.7277 $\pm$ 0.0031 1.2072 $\pm$ 0.0047 5.23 $\pm$ 0.0012 4.67 $\pm$ 0.0057
unmitigated 0.7362 $\pm$ 0.0031 1.2236 $\pm$ 0.0047 5.22 $\pm$ 0.0057 4.68 $\pm$ 0.0015
$(E_{1})$
Analytical value 3.17 2.24
qasm_simulator 3.17 $\pm$ 0.0001 2.24 $\pm$ 0.0047
ibm_fez error mitigated 3.17 $\pm$ 0.0011 2.24 $\pm$ 0.0027
unmitigated 3.16 $\pm$ 0.0061 2.25 $\pm$ 0.0017
ibm_torino error mitigated 3.17 $\pm$ 0.0051 2.24 $\pm$ 0.0027
unmitigated 3.17 $\pm$ 0.0091 2.25 $\pm$ 0.0022
ibm_kyiv error mitigated 3.17 $\pm$ 0.0001 2.24 $\pm$ 0.0047
unmitigated 3.18 $\pm$ 0.00022 2.24 $\pm$ 0.0045
$(E_{2})$
Analytical value -0.0180 -0.0726 -4.45 -3.65
qasm_simulator -0.1179 $\pm$ 0.0069 -0.1433 $\pm$ 0.0054 -4.45 $\pm$ 0.0069 -3.65 $\pm$ 0.0054
ibm_fez error mitigated 0.0715 $\pm$ 0.0050 -0.1080 $\pm$ 0.0073 -4.45 $\pm$ 0.0049 -3.65 $\pm$ 0.0034
unmitigated -0.0415 $\pm$ 0.0036 -0.1120 $\pm$ 0.0040 -4.46 $\pm$ 0.0019 -3.64 $\pm$ 0.0022
ibm_torino error mitigated -0.0718 $\pm$ 0.0112 -0.1219 $\pm$ 0.0112 -4.45 $\pm$ 0.0026 -3.65 $\pm$ 0.0065
unmitigated -0.0718 $\pm$ 0.0112 -0.1219 $\pm$ 0.0112 -4.45 $\pm$ 0.0011 -3.65 $\pm$ 0.0097
ibm_kyiv error mitigated -0.0973 $\pm$ 0.0038 -0.1483 $\pm$ 0.0045 -4.45 $\pm$ 0.0023 -3.65 $\pm$ 0.0054
unmitigated -0.1098 $\pm$ 0.0038 -0.1738 $\pm$ 0.0045 -4.45 $\pm$ 0.0029 -3.66 $\pm$ 0.0010
TABLE I: Simulation results for different backends and methods compared to analytical values
It was established earlier in [34] that we have observed for all parameter $(k,h)$ combinations, negative $\langle V\rangle$ and it is true for all type of quantum computers of IBM.Previously which was proved for 2-qubit system in [34], is also true for our 3-qubit system. Bob can extract greater energy if only $V_{0,2}$ and $V_{1,2}$is observed, since $\langle H_{2}\rangle$ is always positive (Fig. 5). For practical purposes using minimal model was enough as said by the author of the correspond paper, in addition, our model performs even better, which in turn takes QET one step forward. Either way, we have to keep in mind that Bob’s energy becomes smaller when he observes $H_{2}$.
IV Measurements and Quantum Gates
Here we give a comprehensive explanation of the fundamental concepts related to quantum gates and measurement. We utilize the following one-qubit operators with their corresponding matrix representations
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\\ Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}.$ (34)
For the computational basis states, we can use $|0\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},|1\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$, they are eigenstates of $Z:Z|0\rangle=|0\rangle,Z|1\rangle=-|1\rangle$. Another basis vectors are also beign used |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)$. They are eigenstates of $X:X|\pm\rangle=\pm|\pm\rangle,X|+\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)$.
[…]
The rotation of $X,Y,Z$ can be defined by
$R_{X}(\phi)=e^{-i\frac{\phi}{2}X},R_{Y}(\phi)=e^{-i\frac{\phi}{2}Y},R_{Z}(\phi)=e^{-i\frac{\phi}{2}Z}.$ (35)
Two-qubit gate operations were used previously which is also followed by our method adding one extra qubit. In general, a controlled $U$ operation $\Lambda(U)$ is defined by
$\Lambda(U)=|0\rangle|0\rangle\langle 0|\langle 0|\otimes I+|1\rangle|1\rangle\langle 1|\langle 1|\otimes U$ (36)
and the corresponding diagram is drawn as later
By using similar formula of $CNOT(a|0\rangle+b|1\rangle)|0\rangle=a|00\rangle+b|11\rangle$, it is also evident that {strip}
$CCNOT(R_{Y}(\theta_{1}\otimes I\otimes I)|000\rangle=0|000\rangle+a|001\rangle+b|010\rangle+0|011\rangle+c|100\rangle+0|101\rangle+0|110\rangle+d|111\rangle.$ (37)
$CNOT(R_{Y}(I\otimes\theta_{2}\otimes I)|000\rangle=0|000\rangle+a|001\rangle+b|010\rangle+0|011\rangle+c|100\rangle+0|101\rangle+0|110\rangle+d|111\rangle.$ (38)
We can measure $Z_{2}$ and $X_{i}X_{j}$.For $Z_{2}$ the circuit is
and for $V$ the circuit is
We get measurement output as bit string $b_{0}b_{1}b_{2}\in\{000,001,010,011,100,101,110,111\}$. As we know are $-1,1$ are the eigenvalues of $Z$, we can write the bit strings as $1-2b_{2}$. Let $n_{shots}$ denotes the number of time the circuit is repeated, and $count_{b_{0}b_{1}b_{2}}$ be how may times we detect $b_{0}$, $b_{1}$ and $b_{2}$. Therefore $P_{b_{0}b_{1}b_{2}}$ is the probability that a bit string $b_{0}b_{1}b_{2}$ is obtained. Then the expectation value of $Z_{2}$ is computed by the formula
$\langle Z_{2}\rangle=\sum_{b_{0},b_{1}b_{2}}(1-2b_{2})\frac{count_{b_{0}b_{1}b_{2}}}{n_{shots}}.$ (39)
Measurement of $X_{i}X_{j}$ is done by the following circuit
We know that, applying $H$ maps $|0\rangle,|1\rangle$ to the eigenvectors of $X$, |+,|ketket|$\rangle$,|$\rangle$| + ⟩ , | - ⟩ . Once afain, we get the outputs as a bit string of $b_{0}b_{1}b_{2}\in\{000,001,010,011,100,101,110,111\}$. Which we can use to convert into the eigenvalues of $X_{i}X_{j}$ by writing them as $(1-2b_{i})(1-2b_{j})$. Then we get the expectation value of $X_{i}X_{j}$
$\langle X_{i}X_{j}\rangle=\sum_{b_{i},b_{j}}(1-2b_{i})(1-2b_{j})\frac{count_{b_{0}b_{1}b_{2}}}{n_{shots}}.$ (40)
V Some details of the model
This section is comprised of a comprehensive description of the model utilized in our study. Additional information are available in Hotta’s original papers. It is crucial to acknowledge that the lowest energy state of the whole Hamiltonian is not the lowest energy state of local operators.
In order to comprehend the non-triviality of the QET protocol, it is crucial to acknowledge that regardless of the specific unitary operation W1 applied to Bob’s qubit following Alice’s measurement, it is impossible to extract any energy. This can be verified by the equation
$\text{Tr}[\rho_{W}H_{tot}]-\langle E_{0}\rangle=\langle g|W_{1}^{\dagger}H_{tot}W_{1}|g\rangle\geq 0,$ (41)
where
$\rho_{W}=W_{1}^{\dagger}\left(\sum_{\lambda=\pm 1}P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})\right)W_{1}.$ (22)
Bob’s local system’s time evolution, if he does not perform any operations on his own system after Alice’s measurement, is depicted as
$\langle H_{2}(t)\rangle=\text{Tr}[\rho_{M}e^{itH}H_{2}e^{-itH}]$ (42)
$\langle V(t)\rangle=\text{Tr}[\rho_{M}e^{itH}Ve^{-itH}]=0,$ (43)
where $\rho_{M}=\sum_{\lambda=\pm 1}P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})$.
The time evolution of the system, results in energy transfer to Bob’s local system. However, this transfer is simply the propagation of energy in the usual manner. In Quantum Energy Teleportation (QET), energy is not acquired through the natural progression of time inside the system, but rather immediately through communication. Given that we are examining a non-relativistic quantum many-body system, the rate at which energy travels is significantly slower than the speed of light. Optical communication, a kind of classical communication, may transmit information to distant locations at a significantly faster rate than the temporal progression of physical systems. As, QET is recognized as a rapid energy transmission protocol, our method employ efficiency on top of that.
The change in entropy before and after the measurement is
$\Delta S_{AB}=S_{AB}-\sum_{\lambda=\pm 1}p_{\mu}S_{AB}(\mu)$ (44)
VI IMPLICATIONS FOR OUR REAL WORLD
Our findings have significance for the development of novel quantum communication technologies in many timeframes. It is crucial to acknowledge that, similar to quantum teleportation, energy can exclusively be teleported through Local Operations and Classical Communication (LOCC). Implementing the extended QET model we utilized in our demonstration within small facility can be feasibly accomplished in the near future using existing quantum computing and communication capabilities in an efficient manner. A quantum gadget consisting of 3 qubits and a gate depth of 9 can be prepared for immediate experimentation. Anticipated outcomes include advancements in the utilization of quantum memory [19, 20, 21, 22]. Moreover, the crucial task for future applications was to validate QET in many quantum systems and materials that go beyond the minimum model, which we have addressed in this paper.
New concepts like Quantum Oblivious Transfer (QOT), Quantum Block Chain and Quantum Interactive Protocol can be implemented on much user friendly level owing to this newly introduced enhanced QET protocol which in turn increases the efficiency of the existing protocol as well.
Unlimited distance quantum energy teleportation is also available [22]. The capacity to transmit quantum energy across extensive distances will initiate a groundbreaking transformation in quantum communication technology. Put simply, it is possible to create a future where physical quantities can be sent freely and instantly to distant locations over a large-scale Quantum Internet (Network) that is also scalable for mass end level users. a long-distance quantum network already exists in Long Island, New York [23]. Several quantum networks have been established [24, 25, 26]. Implementing (QET) on a quantum level network is anticipated to be feasible by the 2030s, would represent a significant achievement in the progression towards establishing QET on a global scale.
The implementation of a extended long-range Quantum Entanglement Teleportation (QET) will have significant ramifications that exceeds beyond the advancement of information and communication technologies and quantum physics. Information and energy possess physical properties and also have economic implications. The ability to sell physical quantities directly on the quantum network will result in the emergence of a new economic market [27]. Quantum teleportation is a well-established technique that is now being explored for practical applications. Furthermore, the real implementation of QET will grant us access to a multitude of quantum resources. The Hermite operator’s predicted value is commonly referred to as energy, although its usage is not limited to literal energy calculations. Teleported energy can be utilized as a source of energy, as well as for several other purposes. The utilization of quantum entanglement to transfer energy, a cheap physical commodity, has increased its worth. In a quantum marketplace with Alice, Bob, and Charlie, if Bob can obtain a greater amount of energy from Charlie compared to Alice, Bob may choose to engage in transactions with Charlie instead of Alice. Additionally, Bob may have a preference for an entangled state with Charlie. Nevertheless, Bob might opt for Alice based on transaction costs. Many game-theoretic situations can be generated [28, 29, 30, 31, 32]. This suggests that the concept of quantum information economics, which is currently nonexistent, will gain significance in the future.
VII Conclusion
The ground state of a many-body quantum system is usually entangled which can be subjected to various interesting protocols and applications. Although the measurement of a subsystem destroys the entanglement, some energies are injected into the local system, some of which can be retrieved using local operations and classical communication. The minimal Quantum Energy Teleportation (QET) model exploits this phenomenon for two-qubit systems. This paper extends the minimal QET circuit for a 3-qubit system and utilizes squeezed vacuum states with local vacuum regions between three contemporary protocol users. This overcomes the limitation of the low-energy extraction problem and the simulation obtains a retrieval efficiency of 46.4% in contrast with the 11.4% efficiency obtained in the 2-qubit protocol. This result makes the proposed protocol more efficient with applications in quantum energy teleportation and condensed matter physics.
VIII Acknowledgement
We thank Mr. Sabir Md Sanaullah and Mr. Md Zubair for fruitful communication and initial discussions. We acknowledge the use of IBM quantum computers and Quantum Environment. M.R.C. Mahdy acknowledges the support of NSU internal grant and CTRGC grant 2023-24.
IX Competing interests
The author declares that there is no competing financial interests.